# Graph RAG for Legal Norms: A Hierarchical and Temporal Approach

### v1

@@ Abstract
This article proposes an adaptation of Graph Retrieval Augmented Generation (Graph RAG) specifically designed for the analysis and comprehension of legal norms, which are characterized by their predefined hierarchical structure, extensive network of internal and external references and multiple temporal versions. By combining structured knowledge graphs with contextually enriched text segments, Graph RAG offers a promising solution to address the inherent complexity and vast volume of legal data. The integration of hierarchical structure and temporal evolution into knowledge graphs - along with the concept of comprehensive Text Units - facilitates the construction of richer, interconnected representations of legal knowledge. Through a detailed analysis of Graph RAG and its application to legal norm datasets, this article aims to significantly advance the field of Artificial Intelligence applied to Law, creating opportunities for more effective systems in legal research, legislative analysis, and decision support.
@@ 1 Introduction
Legal frameworks are inherently hierarchical, with constitutions, statutes, regulations, and jurisprudence organized in a structured manner. Understanding the relationships between the different levels of this hierarchy is essential for the interpretation and application of Law. Higher-level legal principles (e.g., constitutional provisions) frequently influence or restrict lower-level norms (e.g., specific regulations) – regarding the hierarchy of norms and legal principles, see [1].
Additionally, the dynamic nature of legal norms, evolving through amendments, repeals, and new interpretations, results in multiple temporal versions. The ability to track and understand this evolution [2] is, therefore, essential for determining the applicable legislation at a given moment. These defining characteristics of legal norms, their structural hierarchy (alongside the cross-references) and their temporal dimension, although representing significant analytical challenges, are indispensable for correct application and normative interpretation. Consequently, they demand analysis approaches that explicitly capture both dimensions – precisely the objective of the adapted Graph-based Retrieval-Augmented Generation (Graph RAG) methodology in this article.
@@
This paper is structured as follows. Section 2 first defines the core structural entities – Norm, Component, and Version – essential for modeling legal documents, discussing their hierarchical relationships and the challenges of representing temporal changes. Section 3 then elaborates on our proposed adaptations to the Graph RAG framework, introducing the concept of cumulative Text Units, explaining the rationale for modeling versions as aggregations rather than simple compositions, detailing the inclusion of essential metadata and alteration actions, and discussing strategies for structure-aware retrieval. Lastly, Section 4 concludes the paper by highlighting the key contributions and suggesting avenues for future investigation.
@@ 2 Structural Entities in the Graph
The initial step in both Graph RAG and traditional RAG approaches is to divide each text in the input dataset into segments, or “text chunks”. These segments are then converted into embeddings [7] and stored in a vector database [8]. Various segmentation strategies can be employed, such as sentence segmentation, paragraph segmentation, window-based splitting (according to the embedding model’s window size), and semantic segmentation (dividing text according to its semantic blocks), among others.
The second step involves extracting Entities, along with their properties and relationships, from each text segment generated in the previous step. These extracted entities serve as the “building blocks” or nodes of the knowledge graph, with edges interconnecting them to represent the relationships derived from the normative texts.
@@
    Content Entities
@@
Considering the first group, we can find traditional entities in the text of a legal norm such as types: Person, Organization, Location, Event, Concept (metamodel entity [^3] Legal metamodel, in this context, refers to a conceptual model that defines the fundamental types of entities and relationships in the domain of Law, such as abstract legal concepts, principles, values, etc. Examples would be: ”Right to Life”, ”Principle of Legality”, ”Social Contract”.), Legal Norm (normative reference [^4] Normative reference represents a remission or citation to another legal norm, either in its entirety or to a specific provision.) etc.
Regarding the second group, we find entities that represent the hierarchical structure of the text of a legal norm (its articulation), according to the formation rules of this type of text, and those that represent the versions of this structure over time.
@@ 2.1 Legal Norm Structure
This hierarchical organization not only facilitates the systematic interpretation of legal texts but also enables precise navigation and retrieval of specific legal provisions. The basic unit of a Brazilian legal norm’s articulation is the Artigo (Article). Each Article can be further subdivided into additional provisions, as demonstrated in Figure 1.
Figure 1: Diagram of the hierarchical structure of possible subdivisions for the Article (Artigo) provision in Brazilian legislation. (Diagram shows: Artigo (Article) $\to$ Caput; Artigo (Article) $\to$ Parágrafo (Paragraph); Caput, Parágrafo (Paragraph) $\to$ Inciso (Item); Inciso (Item) $\to$ Alínea (Subitem); Alínea (Subitem) $\to$ Item (Sub-subitem))
@@
Figure 2: Diagram of the hierarchical structure of grouping for the "Article" (Artigo) provision in Brazilian legislation (Diagram shows hierarchy: Articulação (Articulation) $\to$ Parte (Part) $\to$ Livro (Book) $\to$ Título (Title) $\to$ Capítulo (Chapter) $\to$ Seção (Section) $\to$ Subseção (Subsection) $\to$ Artigo (Article))
@@ 2.2 Legal Norm Text Segmentation
Ideally, semantic segmentation [^7] Semantic segmentation is a division of the document based on the meaning and structure of the text, rather than merely on superficial textual criteria such as sentences or paragraphs. should be employed to divide the document into text segments that correspond directly to its structural elements, as illustrated in Figure 3. In this approach, the initial segmentation step would simultaneously extract the structural entities, thereby capturing the hierarchical organization from the outset. The subsequent step would focus on extracting content entities, as outlined in the original Graph RAG proposal.
@@
As an example, a passage of the the “Constituição Federal Brasileira de 1988” (Federal Constitution of Brazil (1988)) received the semantic segmentation (like showed in Figure 3), producing the entity of type Norm, represented by a circle in green, and those of type “Component”, represented in blue in the hierarchy presented in Figure 4.
Figure 4: Illustration of hierarchical semantic segmentation applied to a passage of the Federal Constitution of Brazil (1988), representing the "Norm" (green), "Components" (blue) and their associated texts
@@ 2.3 Text Versioning over Time
Version entities represent the view over time of a norm/component, which can be extracted from its original text, but also from the “compilação multivigente999“Compilação” (Compilation) is a process in which, starting from the original version of a legal norm, all modifications commanded by subsequent norms are applied to its text, up to the specified date, thus producing an unofficial updated text of this norm on this date. It is ”Multivigente” (Multi-temporal) when all dates on which the norm underwent alteration are considered in the Compilation, producing an unofficial updated text (version) for each identified date. (multi-temporal compilation) of its text.
@@
Since the text of a norm is constructed by hierarchically grouping the texts of its individual components, the creation of a new version for any component automatically generates a new version for the entire norm as well as for all aggregating elements/components that include that component.
For example, Figure 5 illustrates the Version entities for the provision “Art. 6” of the "Constituição Federal brasileira” (Brazilian Federal Constitution) [^10] The structuring of the ”Constituição Federal Brasileira de 1988” (Brazilian Federal Constitution of 1988) into ”Norm”, ”Components”, and ”Versions” can be visualized at https://normas.leg.br/?urn=urn:lex:br:federal:constituicao:1988-10-05;1988. Accessed on 15.03.2025., demonstrating how different versions reflect changes over time.
Figure 5: Structure of the provision "Art. 6" of the "Constituição Federal Brasileira" (Brazilian Federal Constitution), illustrating the Version entities (light gray circles) and their respective events and contents over time. The versions represent the different wordings of "Art. 6" on different dates, due to constitutional amendments.
@@ 2.4 Building the Graph
After semantically segmenting the text of a legal norm into a hierarchical structure (tree), as per the proposal presented above, the elements of the hierarchy are then inserted as nodes of the types Norm (with its associated properties, such as URN - Uniform Resource Name, name, signature date etc.) and “Component” (with its properties such as URN, inclusion date etc.) into the Knowledge Graph, as illustrated in Figure 4, which demonstrates the semantic segmentation of a Constitution segment.
Next, for each component, a node of the type Version will be inserted (and its properties such as URN, version date, label/name, text etc.), with its associated text segment [11], as presented in Figure 6 – this associated text segment (text chunk) is fundamental for the Retrieval Augmented Generation process, as it represents the textual content of that specific version that will be retrieved and used as context for the responses generated by the LLM.
@@
For example, the Norm [^11] For a more precise representation in the Graph, specialized entity types can be created, derived from Norm, for each type of norm in the treated legal system. In Brazil, there are, for example: ”Constituição Federal” (Federal Constitution), “Lei Ordinária” (Ordinary Law) etc. Thus, the entity representing a legal norm could be included in the graph using the specialized type. node/entity “Constituição Federal de 1988” (Federal Constitution of 1988), identified by urn [^12] The formation of the urns presented in this work follows the standard proposed by the Brazilian LexML project.
urn:lex:br:federal:constituicao:1988-10-05;1988, has a Dispositivo [^13] Dispositivo” (Provision) is a subtype of Component. For an even more precise representation in the Graph, specialized entity types can be created, derived from Dispositivo, for each type of provision found in the treated legal system. For example, in the Brazilian case, there are types like ”Artigo” (Article), ”Inciso” (Item), “Alínea” (Subitem), “Título” (Title), “Capítulo” (Chapter), presented in figures 1 and 2. Thus, the entity representing a component could be included in the graph using the specialized type. (“Provision”) node/entity “caput do Art. 6º” (caput is the main text of the article), identified by
urn:lex:br:federal:constituicao:1988-10-05;1988!art6_cpt, which has associated the original version Version [^14] For a more precise representation in the Graph, specialized entity types can be created, derived from Version, for each type of version found in the treated legal system. For example, in Brazil, there are types like ”Original”, “Alteração de Texto” (Text Change), “Renumeração” (Renumbering), “Revogação” (Repeal) etc. Thus, the entity representing the version could be included in the graph using the specialized type. node/entity, identified by
urn:lex:br:federal:constituicao:1988-10-05;1988@1988-10-05!art6_cpt, that represents the version on the date 1988-10-05 (original version) of the text of the caput of Art. 6 of Federal Constitution of 1988. For this Version node/entity, we have the associated Text chunk with the text bellow, as illustrated in Figure 5:
@@
Unlike other fields, where the text remains static over time (for example, the transcribed speech of a Parliamentarian in a Congress session will not change over time), the text over time of a legal norm, or of a provision, can be represented by a set of Versions, as exemplified by Figure 7. In this figure, four versions of the caput of Art. 6, found in Figure 5, are displayed: one original and three resulting from changes promoted by “Emendas Constitucionais” (Constitutional Amendments).
@@
The proposal of this work is not only to incorporate the entities of the hierarchical structure of the legal norm’s text into the knowledge graph, but also to include the Version entities of each component of the norm. It is important to note that the text segments (text chunks), identified by the original text and by the versions of the text over time of the norm, will be linked to the Version entities, and not to Component entities, as they represent the text of the component at different moments in time, as shown in Figure 7 - in practice, this link could be materialized as a property of the Version entity in the graph, just like a property for the embedding of its content.
@@ 3.1 Text Units
Regarding the text segment (text chunk) associated with each Version entity, in cases of subdivisions of a caput, such as “parágrafos” (paragraphs) and enumeration elements (“incisos” (items) and “alíneas” (subitems)), we suggest adopting the strategy proposed by [12]. In this strategy, the text segment to be considered for a component is not only the text directly resulting from the semantic segmentation of the norm (or its compilation over time) but the semantically complete text of the component, derived from post-segmentation (and post-compilation) processing. The cited author recognizes the hierarchical and interdependent nature of the components of a legal norm, where, for example, the meaning of an enumeration component (e.g., “inciso” (item)) is frequently dependent on the context of its parent component (e.g., the caput).
@@
but rather the text formed by the content of the article’s caput (on the same date) plus the content of the item itself, which can also be visualized in Figure 8:
@@
The version of an article’s caput will have as linked text its own text accumulated with the text from the tree of its child components, as exemplified in Figure 9. The same applies to the article, the groupers and the norm (as a whole), accumulating its own text with the text from their respective trees. According to the cited author, this would bring more “aboutness” [^15] The term ”aboutness” refers to the capacity of a text to faithfully represent its central theme. For theoretical discussions, see [17]. to the texts linked to the entities [13]. In other words, the cumulative “Text Unit”, by including the textual context of the parent, or child, components, represents the central theme and the legal meaning of the component in question more completely and comprehensively, increasing how well the text captures "what it is about" the provision.
@@
Considering this new approach, we will call the texts associated with the entities Text Units, instead of Text Chunks, as they are not necessarily the pieces of text resulting directly from the segmentation of the norms’ texts, but rather the contextually complete content of each component.

### v2

@@ Abstract
This article proposes an adaptation of Graph Retrieval Augmented Generation (Graph RAG) specifically designed for the analysis and comprehension of legal norms, which are characterized by their predefined hierarchical structure, extensive network of internal and external references and multiple temporal versions. By combining structured knowledge graphs with contextually enriched text segments, Graph RAG offers a promising solution to address the inherent complexity and vast volume of legal data. The integration of hierarchical structure and temporal evolution into knowledge graphs - along with the concept of comprehensive Text Units - facilitates the construction of richer, interconnected representations of legal knowledge. Through a detailed analysis of Graph RAG and its application to legal norm datasets, this article aims to advance the field of Artificial Intelligence applied to Law, creating opportunities for more effective systems in legal research, legislative analysis, and decision support.
@@ 1 Introduction
Legal frameworks are inherently hierarchical. This structure manifests both between different types of norms (constitutions, statutes, regulations) and within individual legal texts themselves. Higher-level legal principles (e.g., constitutional provisions) frequently influence or restrict lower-level norms (e.g., specific regulations) – regarding the hierarchy of norms and legal principles, see [1]. On the other hand, legal norms are typically organized internally through grouping provisions such as titles, chapters, sections, and further subdivided into specific provisions like articles, paragraphs, items, clauses etc. Understanding the relationships across all levels of this hierarchy – both inter-norm and intra-norm – is essential for the interpretation and application of Law.
Additionally, the dynamic nature of legal norms, evolving through amendments, repeals, and new interpretations, results in multiple temporal versions. The ability to track and understand this evolution [2] is, therefore, essential for determining the applicable legislation at a given moment.
These defining characteristics of legal norms, although representing significant analytical challenges, are indispensable for correct application and normative interpretation. Consequently, they demand analysis approaches that explicitly capture the structural and temporal dimensions of the legal norms, precisely the objective of the adapted Graph-based Retrieval-Augmented Generation (Graph RAG) methodology in this article.
@@
This paper is structured as follows. Section 2 first defines the core structural entities – Norm, Component, and Version – essential for modeling legal documents, discussing their hierarchical relationships and the challenges of representing temporal changes. Section 3 then elaborates on our proposed adaptations to the Graph RAG framework, introducing the concept of Text Units, explaining the rationale for modeling versions as aggregations rather than simple compositions, detailing the inclusion of essential metadata and alteration actions, and discussing strategies for structure-aware retrieval. Lastly, Section 4 concludes the paper by highlighting the key contributions and suggesting avenues for future investigation.
@@ 2 Structural Entities in the Graph
The initial step in both Graph RAG and traditional RAG approaches is to divide each text in the input dataset into segments, or “text chunks” - these segments are, throughout the process, converted into embeddings [7] and stored in a vector database [8]. Various segmentation strategies can be employed, such as sentence segmentation, paragraph segmentation, window-based splitting (according to the embedding model’s window size), and semantic segmentation (dividing text according to its semantic blocks), among others.
The second step in Graph RAG involves extracting Entities, along with their properties and relationships, from each text segment generated in the previous step. These extracted entities serve as the “building blocks” or nodes of the knowledge graph, with edges interconnecting them to represent the relationships derived from the normative texts.
@@
    Content Entities: traditional entities in the text of a legal norm such as types Person, Organization, Location, Event, Concept (metamodel entity [^3] Legal metamodel, in this context, refers to a conceptual model that defines the fundamental types of entities and relationships in the domain of Law, such as abstract legal concepts, principles, values, etc. Examples would be: ”Right to Life”, ”Principle of Legality”, ”Social Contract”.), Legal Norm (normative reference [^4] Normative reference represents a remission or citation to another legal norm, either in its entirety or to a specific provision.) etc.
@@
    • entities that represent the hierarchical structure of the text of a legal norm (its articulation), according to the formation rules of this type of text;
    • entities that represent the versions of this structure over time.
@@ 2.1 Legal Norm Structure
This hierarchical organization not only facilitates the systematic interpretation of legal texts but also enables precise navigation and retrieval of specific legal provisions.
The basic unit of a Brazilian legal norm’s articulation is the Artigo (Article). Each Article can be further subdivided into additional provisions, as demonstrated in Figure 1.
Figure 1: Diagram of the hierarchical structure of possible subdivisions for the Article (Artigo) provision in Brazilian legislation. It shows Artigo (Article) $\to$ Caput (main section); Artigo (Article) $\to$ Parágrafo (Paragraph); Caput, Parágrafo (Paragraph) $\to$ Inciso (Item); Inciso (Item) $\to$ Alínea (Subitem); Alínea (Subitem) $\to$ Item (Sub-subitem).
@@
Figure 2: Diagram of the hierarchical structure of grouping for the "Article" (Artigo) provision in Brazilian legislation. It shows Articulação (Articulation) $\to$ Parte (Part) $\to$ Livro (Book) $\to$ Título (Title) $\to$ Capítulo (Chapter) $\to$ Seção (Section) $\to$ Subseção (Subsection) $\to$ Artigo (Article).
@@ 2.2 Legal Norm Text Segmentation
Ideally, semantic segmentation [^7] Semantic segmentation is a division of the document based on the meaning and structure of the text, rather than merely on superficial textual criteria such as sentences or paragraphs. should be employed to divide the document into text segments that correspond directly to its structural elements, as illustrated in Figure 3. In this approach, the initial segmentation step would simultaneously extract the structural entities, thereby capturing the hierarchical organization from the outset - the subsequent step would focus on extracting content entities, as outlined in the original Graph RAG proposal.
@@
As an example, a passage of the the “Constituição Federal Brasileira de 1988” (Federal Constitution of Brazil (1988)) received the semantic segmentation (like showed in Figure 3), producing the entity of type Norm, represented by a circle in green, and those of type Component, represented in blue in the hierarchy presented in Figure 4.
For a more precise representation in the Graph, specialized entity types, derived from Norm, can be created for each type of norm in the treated legal system. In Brazil, there are, for example: "Constituição Federal" (Federal Constitution), “Lei Ordinária" (Ordinary Law) etc. Thus, the entity representing a legal norm could be included in the graph using the specialized type.
Specialized entity types, derived from Component, can also be created for each type of provision found in the treated legal system. For example, in the Brazilian case, there are types like "Artigo" (Article), "Inciso” (Item), “Alínea" (Subitem), “Título” (Title), “Capítulo” (Chapter), presented in figures 1 and 2. Thus, the entity representing a component could be included in the graph using the specialized type.
Figure 4: Illustration of hierarchical semantic segmentation, and typification of the structural entities, applied to a passage of the Federal Constitution of Brazil (1988), representing the "Norm" (green), "Components" (blue) and their associated texts.
@@ 2.3 Text Versioning over Time
Version entities represent the view over time of a norm/component, which can be extracted from its original text, but also from the “compilação multivigente999“Compilação” (Compilation) in Brazil is a process in which, starting from the original version of a legal norm, all modifications commanded by subsequent norms are applied to its text, up to the specified date, thus producing an unofficial updated text of this norm on this date. It is ”Multivigente” (Multi-temporal) when all dates on which the norm underwent alteration are considered in the Compilation, producing an unofficial updated text (version) for each identified date. (multi-temporal compilation) of its text.
@@
Since the text of a norm is constructed by hierarchically grouping the texts of its individual components, the creation of a new version for any component automatically generates a new version for the entire norm as well as for all grouping elements/components that include that component.
For example, Figure 5 illustrates the Version entities for the provision “Art. 6” (and its caput - main section) of ’Constituição Federal brasileira’ (Brazilian Federal Constitution)" [^10] The structuring of the ”Constituição Federal Brasileira de 1988” (Brazilian Federal Constitution of 1988) into ”Norm”, ”Components”, and ”Versions” can be visualized at https://normas.leg.br/?urn=urn:lex:br:federal:constituicao:1988-10-05;1988. Accessed on 15.03.2025., demonstrating how different versions reflect changes over time.
Figure 5: Structure of the provision "Art. 6" (and its caput - main section) of Brazilian Federal Constitution, illustrating four Version entities over time (light gray circles) and their respective generated events and resulting contents. The versions represent the different wordings of the caput on different dates, due to constitutional amendments.
@@ 2.4 Building the Graph
After semantically segmenting the text of a legal norm into a hierarchical structure (tree), as per the proposal presented above, the elements of the hierarchy are then inserted as nodes of the types Norm (with its associated properties, such as URN - Uniform Resource Name, name, signature date etc.) and Component (with its properties such as URN, optional inclusion date etc.) into the Knowledge Graph, as illustrated in Figure 4, which demonstrates the semantic segmentation of a Constitution passage.
Next, for each component, a node of the type Version will be inserted with its associated text segment [11] (and its properties such as URN, version date, label/name, text etc.) as presented in Figure 6 – this associated text segment (text chunk) is fundamental for the Retrieval Augmented Generation process, as it represents the textual content of that specific version that will be retrieved and used as context for the responses generated by the LLM.
@@
For example, the Norm node/entity “Constituição Federal de 1988” (Federal Constitution of 1988)
identified by urn [^11] The formation of the urns presented in this work follows the standard proposed by the Brazilian LexML project. Available at: https://projeto.lexml.gov.br. Accessed on: 15 Mar 2025.
has a Component node/entity “caput do Art. 6º” (caput is the main text of the article) identified by
which has associated the original version Version node/entity, identified by
that represents the version on the date 1988-10-05 (original version) of the text of the caput of Art. 6 of Federal Constitution of 1988. For this Version node/entity, we have a associated Text chunk with the text bellow - as illustrated in Figure 5:
@@
("Social rights include education, health, work, leisure, security, social security, protection of motherhood and childhood, and assistance to the destitute, as prescribed by this Constitution.")
Unlike other fields, where the text remains static over time (for example, the transcribed speech of a Parliamentarian in a Congress session will not change over time), the text over time of a legal norm, or of a component, can be represented by a set of Versions.
In Figure 7, four versions of the caput of Art. 6, found in Figure 5, are displayed: one original and three resulting from changes (text change) promoted by “Emendas Constitucionais” (Constitutional Amendments) - for a more precise representation in the Graph, specialized entity types, derived from Version, can be created for each type of version found in the treated legal system. For example, in Brazil, there are types like "Original Version”, “Alteração de Texto” (Text Change Version), “Renumeração” (Renumbering Version), “Revogação” (Repeal Version) etc. Thus, the entity representing the version could be included in the graph using the specialized type.
@@
The proposal of this work is not only to incorporate the entities of the hierarchical structure of the legal norm’s text into the knowledge graph, but also to include the Version entities of each component of the norm. It is important to note that the text segments (text chunks), identified by the original text and by the versions of the text over time of the norm, will be linked to the Version entities, and not to Component entities, as they represent the content of the component at different moments in time, as shown in Figure 7 - in practice, this link could be materialized as a property of the Version entity in the graph, just like a property for the embedding of its content.
@@ 3.1 Text Units
Regarding the text segment (text chunk) associated with each Version entity, in cases of subdivisions of a caput, such as “parágrafos” (paragraphs) and enumeration elements (“incisos” (items) and “alíneas” (subitems)), we suggest adopting the strategy proposed by [12]. In this strategy, the text segment to be considered for a component is not only the text directly resulting from the semantic segmentation of the norm (or from its compilation over time) but the semantically complete text of the component, derived from post-segmentation (and post-compilation) processing. The cited author recognizes the hierarchical and interdependent nature of the components of a legal norm, where, for example, the meaning of an enumeration component (e.g., “inciso” (item)) is frequently dependent on the context of its parent component (e.g., the caput).
@@
but rather the text formed by the content of the article’s caput (on the same date) plus the content of the item itself, which can also be visualized in Figure 8 - the text of the caput was placed as a prefix in square brackets.
@@
The version of an article’s caput will have as linked text its own text accumulated with the text from the tree of its child components, as exemplified in Figure 9. The same applies to the article, the groupers and the norm (as a whole), accumulating its own text with the text from their respective trees.
According to the cited author, this would bring more “aboutness” [^12] The term ”aboutness” refers to the capacity of a text to faithfully represent its central theme. For theoretical discussions, see [17]. to the texts linked to the entities [13]. In other words, the cumulative “Text Unit”, by including the textual context of the parent, or child, components, represents the central theme and the legal meaning of the component in question more completely and comprehensively, increasing how well the text captures "what it is about" the provision.
@@
Considering this new approach, we will call the texts associated with the entities Text Units, instead of Text Chunks, as they are not necessarily the pieces of text resulting directly from the segmentation (or compilation) of the norms’ texts, but rather the contextually complete content of each component.